Stock Y has a beta of 0.9 and an expected return of 8.67 percent. Stock Z has a beta of 2.2 and an expected return of 14.5 percent. What would the risk-free rate (in percent) have to be for the two stocks to be correctly priced relative to each other? Answer to two decimals. Thank you!!
Answer #1

NN Expected Return = Risk free rate+beta*Risk Premium 0.0867 = Rf+0.9*Rp 0.145 = Rf+2.2*Rp 0.0583 = 1.3Rp = 0.0583/1.3 0.0448